OpenStudy (anonymous):

condense this to a single quantity: 1/2lnX +ln(x+3)-ln(x^2 +1) thanks.

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

This appears to become a quadratic equation

OpenStudy (anonymous):

really? well im not sure because its logarithms and u need to condense it.

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

well it is wrong, so don't write it!

OpenStudy (anonymous):

i made a mistake

OpenStudy (anonymous):

oh i see could you please explain it to me?

OpenStudy (anonymous):

\[\frac{1}{2}\ln(x)=\ln(\sqrt{x})\]

OpenStudy (anonymous):

because \[\ln(x^n)=n\ln(x)\]

OpenStudy (anonymous):

here i am going from the right hand side to the left hand side.

OpenStudy (anonymous):

\[\frac{1}{2}\ln(x)=\ln(x^{\frac{1}{2}})=\ln(\sqrt{x})\]

OpenStudy (anonymous):

then we use \[\ln(ab) = \ln(a)+ln(b)\] again from the right to the left.

OpenStudy (anonymous):

so we get \[\ln(\sqrt{x})+\ln(x+3)=\ln(\sqrt{x}(x+3))\]

OpenStudy (anonymous):

then we use \[\ln(a)-\ln(b)=\ln(\frac{a}{b})\]

OpenStudy (anonymous):

to get \[\ln(\sqrt{x}(x+3))-\ln(x^2+1)=\ln(\frac{\sqrt{x}(x+3)}{x^2+1})\]

OpenStudy (anonymous):

i made a mistake on the first try and wrote \[\frac{1}{2}\ln(x)=\ln(x^2)\] which is wrong

OpenStudy (anonymous):

there now i deleted it. hope the steps are clear

OpenStudy (anonymous):

ohhh no i understand thanks so much!!:]

OpenStudy (anonymous):

welcome!